OpenStudy (anonymous):

solve using quadratic formula. x^2+(x+10)^2=50 solve using quadratic formula. x^2+(x+10)^2=50 @Mathematics

OpenStudy (anonymous):

\[x ^{2}+x ^{2}+20x+100-50=0\] \[2x ^{2}+20x+50=0\] a=2 b=20 c=50 \[\frac{-b \pm \sqrt{b ^{2}-4ac}}{2a}\] \[\frac{-20 \pm \sqrt{20 ^{2}-4*2*50}}{2*2}\]

OpenStudy (anonymous):

\[\frac{-20 \pm \sqrt{400-400}}{4}\]\[\frac{-20 }{4}\]=-5

OpenStudy (anonymous):

x=-5